94, 68, 72, 88, 96, 88, 89, 97 whats the median
Answer by jim_thompson5910(35256)

Original Data Set: 94, 68, 72, 88, 96, 88, 89, 97

Sorted Data Set: 68, 72, 88, 88, 89, 94, 96, 97


Start with the list of numbers shown below

68, 72, 88, 88, 89, 94, 96, 97


There are 8 values. This is an even number of values. So the median will be in between the two middle-most numbers (the median may or may not be in the list -- there are no guarantees).


More specifically, the median will be in between the elements at slot numbers 4 and 5 (divide 8 in half to get 4).


This means you can count 4 places over to get to the value of 88. The next number over is 89


The median will be the average of the two values 88 and 89. So average them to get (88 + 89)/2 = (177)/2 = 88.5


Or you can do it this way

Step 1) Start with the list of numbers shown below

68, 72, 88, 88, 89, 94, 96, 97

---------------------------------------------------------------------

Step 2) Erase the outer pair of numbers (68 on the very left and 97 on the very right) to get the new list:

72, 88, 88, 89, 94, 96

Repeat this until the list reduces to exactly two values.

---------------------------------------------------------------------

Step 3) Erase the outer pair of numbers (72 on the very left and 96 on the very right) to get the new list:

88, 88, 89, 94

---------------------------------------------------------------------

Step 4) Erase the outer pair of numbers (88 on the very left and 94 on the very right) to get the new list:

88, 89

---------------------------------------------------------------------

We're left with exactly two numbers: 88 and 89. So we are done erasing values.


Average the two values to get (88 + 89)/2 = (177)/2 = 88.5


So the median is 88.5